Confused. Help.
Use the change-of-base rule to express log[3]12 in terms of common logarithms, in terms of natural logarithms, and correct to four decimal places.

Found 2 solutions by rothauserc, richard1234:
Answer by rothauserc(4718)

the change of base logarithm rule is
log (base a) of x = log (base b) of x / log (base b) of a
1. common logarithms ( this is base 10)
log base 3 of 12 = log (base 10) of 12 / log (base 10) of 3 = 1.079181246 / 0.47712125472 = 2.261859507 = 2.2619
2. natural logarithm ( this is base e)
log base 3 of 12 = log (base e) of 12 / log (base e) of 3 = 2.4849066498 / 1.0986122887 = 2.261859507 = 2.2619

Answer by richard1234(7193)